OpenStudy (anonymous):

Two trains leave stations 588 miles apart at the same time and travel toward each other. One train travels at 115 miles per hour while the other travels at 95 miles per hour. How long will it take for the two trains to meet?

OpenStudy (uri):

There's a formula for this one. @kdog55

OpenStudy (uri):

I seriously don't remember formula's from my math class. But I'll google.

OpenStudy (anonymous):

Think of this as one train standing still and the other moving toward it at the sum of their speeds until they meet Let t = time in hours until they meet Equation for the moving train: 588=(115+95)t 210t=588 t=2.8hrs

OpenStudy (uri):

It's not letting me post the formula. Hold on.

OpenStudy (anonymous):

2.8 hrs = 2hrs and (0.8 *60) mins = 2hrs and 48 mins

OpenStudy (uri):

http://i.imgur.com/4hVc4Rb.png @kdog55

OpenStudy (anonymous):

what does s and t mean in the formula? @uri

OpenStudy (anonymous):

Two cars leave towns 400 kilometers apart at the same time and travel toward each other. One car's rate is 12 kilometers per hour more than the other's. If they meet in 2 hours, what is the rate of the faster car?

OpenStudy (anonymous):

can u help with this one @sangya21 @bibby @Becki

OpenStudy (uri):

S stands for Speed and T stands for Time and D stands for Distance.

OpenStudy (anonymous):

so how would i put that together

OpenStudy (anonymous):

@uri